# 4D Gaussian Splatting: Modeling Dynamic Scenes with Native 4D Primitives

### v1

@@ Abstract
Dynamic 3D scene representation and novel view synthesis from captured videos are crucial for enabling immersive experiences required by AR/VR and metaverse applications. However, this task is challenging due to the complexity of unconstrained real-world scenes and their temporal dynamics. In this paper, we frame dynamic scenes as a spatio-temporal 4D volume learning problem, offering a native explicit reformulation with minimal assumptions about motion, which serves as a versatile dynamic scene learning framework. Specifically, we represent a target dynamic scene using a collection of 4D Gaussian primitives with explicit geometry and appearance features, dubbed as 4D Gaussian splatting (4DGS). This approach can capture relevant information in space and time by fitting the underlying spatio-temporal volume. Modeling the spacetime as a whole with 4D Gaussians parameterized by anisotropic ellipses that can rotate arbitrarily in space and time, our model can naturally learn view-dependent and time-evolved appearance with 4D spherindrical harmonics. Notably, our 4DGS model is the first solution that supports real-time rendering of high-resolution, photorealistic novel views for complex dynamic scenes. To enhance efficiency, we derive several compact variants that effectively reduce memory footprint and mitigate the risk of overfitting. Extensive experiments validate the superiority of 4DGS in terms of visual quality and efficiency across a range of dynamic scene-related tasks (e.g., novel view synthesis, 4D generation, scene understanding) and scenarios (e.g., single object, indoor scenes, driving environments, synthetic and real data).
Index Terms:
Dynamic scenes, novel view synthesis, Gaussian splatting, neural video synthesis.
I Introduction
Modeling dynamic scenes from 2D images and rendering novel views in real-time is crucial in computer vision and graphics, which has received increasing attention from both industry and academia because of its potential value in a wide range of AR/VR applications. Whilst recent breakthroughs (e.g., NeRF [1]) have achieved photorealistic rendering for static scenes [2, 3], handling dynamic scenes remains a significant challenge. The temporal dynamics of scenes introduce additional complexity. The central challenge lies in preserving intrinsic correlations and sharing relevant information across different time steps, all while minimizing interference between unrelated spacetime locations.
@@ I Introduction
The second group posits that scene dynamics underlie the motion and deformation of a single 3D model [9, 10, 11, 12]. These approaches explicitly model scene motion, potentially improving the utilization of spatial and temporal correlations. However, global tracking of object motion involves inherent ambiguities and risks singularities when dealing with complex motions in real-world scenes, such as sudden appearances and disappearances, which limits their capability and scalability in unconstrained environments.
@@
To address these fundamental limitations, in this work we reformulate the modeling of dynamic scenes as a generic, spatio-temporal 4D volume learning problem. This novel perspective not only introduces an explicit representation model, but also makes minimal assumptions about how motion is composed, paving the way for a versatile dynamic scene learning framework. Concretely, we propose a native 4D Gaussian Splatting (4DGS) model that fully captures all the information involved in the temporal dynamics of a 3D scene. Notably, the 4D Gaussians parameterized by anisotropic ellipses that can rotate arbitrarily in space and time we introduce fits naturally within the intrinsic spatio-temporal manifold. Furthermore, we introduce Spherindrical Harmonics, a generalized form of Spherical Harmonics, to represent time evolution of appearance in dynamic scenes, along with dedicated splatting-based rendering. Our 4DGS model is the first to support end-to-end training and real-time rendering of high-resolution, photorealistic novel views of complex dynamic scenes.
@@
We make the following contributions: (i) We cast dynamic scenes as a spatio-temporal 4D volume learning problem, a generic reformulation with minimal assumptions, and explicit representation. (ii) We propose a native 4DGS model characterized by anisotropic 4D Gaussian primitives equipped with 4D Spherindrical Harmonics that can model the time evolution of view-dependent color in dynamic scenes, as well as a dedicated splatting-based rendering pipeline. (iii) Extensive experiments validate the superiority of 4DGS over previous alternatives in terms of visual quality and efficiency across a variety of dynamic scene-based tasks (e.g., novel view synthesis, content generation, scene understanding) and scenarios (e.g., single object, indoor and outdoor scenes, synthetic and real data).
@@
A preliminary version of this work has been presented [13]. Moving beyond our initial focus on novel view synthesis, here we significantly extend and explore the proposed 4DGS from a generic dynamic scene model perspective. In comparison, we introduce several important improvements: (i) Introducing more compact variants: The 4D volume of dynamic scenes is inherently large, which can lead to rapid increases in the 4DGS size. To address storage demands and reduce memory footprint, we explore model compression across multiple dimensions, including parameter size, the number of Gaussian points, and structural redundancy. We also highlight the benefits of this approach for optimization regularization. (ii) Application extension: We extend our model to handle large-scale dynamic urban scenes, which present additional challenges such as sparse viewpoints and boundless environments. We also address 4D generation from monocular videos with limited observations [14] and enhance 3D scene understanding, including object segmentation across both space and time. (iii) Extensive evaluation: We conduct thorough evaluations of model compression and the newly added applications.
II Related work
II-A Novel view synthesis
Novel view synthesis for static scenes
@@ II-A Novel view synthesis
The field of novel view synthesis has received widespread attention initiated by Neural Radiance Fields (NeRF) [1]. Querying a MLP for hundreds of points along each ray, NeRF is significantly limited in training and rendering. Subsequent works attempted to improve the speed [15, 16, 17, 18, 19, 20] and enhance rendering quality [21, 3, 2, 22, 23]. Later, 3D Gaussian Splatting (3DGS) [24] achieves real-time rendering at high resolution thanks to its fast differentiable rasterization pipeline while retaining the merits of volumetric representations. Subsequent works [25, 26] have improved rendering [27, 28, 29, 30, 31] and geometric accuracy [32, 33, 34]. Here, we explore the potential of Gaussian primitives in representing dynamic scenes.
NeRF-based dynamic view synthesis
@@
Synthesizing novel views of a dynamic scene in a desired time is a more challenging task. The intricacy lies in capturing the intrinsic correlation across different timesteps. Inspired by NeRF, one research line attempts to learn a 6D plenoptic function represented by a well-tailored structure [4, 6, 7, 5, 8]. However, these methods struggle with the coupling between parameters. An alternative approach explicitly models continuous motion or deformation [9, 10, 11, 12]. Among them, point-based approaches have consistently been deemed promising.
Compression of 3DGS
@@
Despite impressive novel view synthesis performance, 3DGS imposes substantial storage requirements due to the associated attributes of a large number of Gaussians. To reduce the memory footprint, [35] uses the hash grid to encode the color and quantifies other parameters through residual vector quantization (RVQ) [36], while utilizing a learned volume mask to eliminate Gaussians with small size. [37] introduces a scale- and resolution-aware redundancy score as a criterion for removing redundant primitives. [38] compresses parameters of part of Gaussians in a sensitivity-aware manner. [39] stores the related Gaussians in the same anchor and dynamically spawns neural Gaussians during rendering. [40] further organizes sparse anchor points using a structured hash grid. We systematically explore these compression approaches with our dynamic scene representations.
Dynamic 3D Gaussians
@@
There have been recent efforts [12, 41, 42, 43, 44, 45, 46, 47] in extending 3DGS for dynamic scenes. A straight approach is to incrementally reconstruct the dynamic scene via frame-by-frame training [12]. [41, 42, 43] model the geometry and dynamics of scenes by joint optimizing Gaussians in canonical space and a deformation field. [44] encourages locality and rigidity between points by factorizing the motion of the scene into a few neural trajectories. These representations hold the topological invariance and low-frequency motion prior, thus well-suited for reconstructing dynamic scenes from monocular videos. However, they assume that dynamic scenes are induced by a fixed set of 3D Gaussians and that all scene elements remain visible. Circumventing the need to maintain ambiguous global tracking, our approach can more flexibly deal with complex real-world scenes.
II-B Applications with 3D/4D scene representations
Driving scene synthesis
@@ II-B Applications with 3D/4D scene representations
Efficient simulation of realistic sensor data in diverse driving scenes is crucial for scaling autonomous driving systems. Early works [48, 49] employ traditional graphics techniques but suffer from domain gap of synthesized data and laborious manual modeling. [50] creates realistic novel driving videos by inserting 3D assets according to the depth map.  [51, 52, 53, 54, 55] utilize neural implicit fields to reconstruct the scene from captured driving data, enabling novel view synthesis. Recent works [56, 57, 58, 59] have applied 3DGS to driving scenes, significantly improving fidelity while achieving real-time simulation. Other notable efforts focus on reconstructing scenes without relying on 2D masks or 3D bounding boxes of dynamic objects. Among them, SUDS [60] and EmerNeRF [61] decompose dynamic and static regions by optimizing optical flow, while $S^{3}$aussian [62] and PVG [63] achieve this through self-supervision.
3D/4D content generation
@@
Generative capabilities of 3D/4D representations have been explored. In 3D generation, DreamFields [64] and DreamFusion [65] utilize NeRF as a base 3D representation. Magic3D [66] incorporates a textured mesh based on the deformable tetrahedral grid in the refinement stage. Also, 3DGS has been applied to 3D generation [67, 68, 69, 70]. For 4D generation, 4D radiance fields [7, 71, 6] have been adopted [72, 71, 14], followed by Gaussian-based representations [41, 42, 13, 46] explored in [73, 74, 75, 76, 77, 78]. We show the advantages of 4DGS in terms of both generation quality and speed in video-to-4D generation task.
3D/4D scene segmentation
@@
Following the success of Segment Anything Model (SAM) [79] in 2D image segmentation, recent works [80, 81, 82, 83, 84, 85, 86, 87] extend NeRF and 3DGS for scene understanding and segmentation. SAGA [81] applies multi-granularity segmentation to 3DGS using a scale gate mechanism. SA4D [80] conducts 4D segmentation based on deformed 3DGS with single-granularity. We validate the general efficacy of 4DGS for 4D scene understanding.
III Method
@@
We propose a generic scene representation, 4D Gaussian splatting (4DGS), for modeling dynamic scenes, as shown in Figure 2. Next, we will delineate each component of 4DGS and corresponding optimization process. In Section III-A, we review 3D Gaussian splatting [24]. In Section III-B, we detail how to represent dynamic scenes and synthesize novel views with our native 4D primitives, i.e., 4D Gaussian. Finally, the optimization framework will be introduced in Section III-C.
III-A Preliminary: 3D Gaussian splatting
@@
Representation of 3D Gaussians
In 3DGS, a scene is represented as a cloud of 3D Gaussians. Each Gaussian has a theoretically infinite scope and its influence on a given spatial position $x\in\mathbb{R}^{3}$ defined by an unnormalized Gaussian function:
@@
Differentiable rasterization via Gaussian splatting
In rendering, given a pixel $(u,v)$ in view $\mathcal{I}$ with extrinsic matrix $E$ and intrinsic matrix $K$, its color $\mathcal{I}(u,v)$ can be computed by blending visible 3D Gaussians that have been sorted according to their depth, as described below:
@@
III-B 4D Gaussian for dynamic scenes
Reformulation
To enable Gaussian splatting for modeling dynamic scenes, reformulation is necessary. In dynamic scenes, a pixel under view $\mathcal{I}$ can no longer be indexed solely by a pair of spatial coordinates $(u,v)$ in the image plane; But an additional timestamp $t$ comes into play and intervenes. Formally this is formulated by extending equation (2) as:
$\mathcal{I}(u,v,t)=\sum^{N}_{i=1}(p_{i}(u,v,t)\alpha_{i}c_{i}(d)\prod^{i-1}_{j=1}(1-p_{j}(u,v,t)\alpha_{j})).$ (5)
@@
$\mathcal{I}(u,v,t)=\sum^{N}_{i=1}(p_{i}(t)p_{i}(u,v|t)\alpha_{i}c_{i}(d)\prod^{i-1}_{j=1}(1-p_{j}(t)p_{j}(u,v|t)\alpha_{j})).$ (6)
@@
Subsequently, it comes to the question of how to represent a 4D Gaussian. An intuitive solution is that we adopt a distinct perspective for space and time. Consider that $(x,y,z)$ and $t$ are independent of each other, i.e., $p_{i}(x,y,z|t)=p_{i}(x,y,z)$, then equation (6) can be implemented by adding an extra 1D Gaussian $p_{i}(t)$ into the original 3D Gaussian. This design can be viewed as imbuing a 3D Gaussian with temporal extension, or weighting down its opacity when the rendering timestep is away from the expectation of $p_{i}(t)$. However, our experiment shows that while this approach can achieve a reasonable fitting of 4D manifold, it is difficult to capture the underlying motion of the scene (see Section VIII-C for details).
Representation of 4D Gaussian
To address the mentioned challenge, we suggest treating the time and space dimensions equally by formulating a coherent integrated 4D Gaussian model. Similar to [24], we parameterize its covariance matrix $\Sigma$ as the configuration of a 4D ellipsoid for easing model optimization:
@@
Proof of unnormalized Gaussian property
The previous formulation treats the unnormalized Gaussian defined in equation (1) as a specialized probability distribution and assumes it can be factorized into the product of the unnormalized Gaussian functions of its conditional and marginal distributions. In this paragraph, we provide a concise proof for this property.
@@
4D spherindrical harmonics
The view-dependent color $c_{i}(d)$ in equation (6) is represented by a series of SH coefficients in the original 3DGS. To more faithfully model the dynamic scenes of the real world, we need to enable appearance variation with both varying viewpoints and evolving time.
@@
where $Y^{m}_{l}$ is the 3D spherical harmonics. The index $l\geq 0$ denotes its degree, and $m$ is the order satisfying $-l\leq m\leq l$. The index $n$ is the order of the Fourier series. The 4D spherindrical harmonics form an orthonormal basis in the spherindrical coordinate system.
III-C Training

### v2

@@ Abstract
Dynamic 3D scene representation and novel view synthesis are crucial for enabling immersive experiences required by AR/VR and metaverse applications. It is a challenging task due to the complexity of unconstrained real-world scenes and their temporal dynamics. In this paper, we reformulate the reconstruction of a time-varying 3D scene as approximating its underlying spatiotemporal 4D volume by optimizing a collection of native 4D primitives, i.e., 4D Gaussians, with explicit geometry and appearance modeling. Equipped with a tailored rendering pipeline, our representation can be end-to-end optimized using only photometric supervision while free viewpoint viewing at interactive frame rate, making it suitable for representing real world scene with complex dynamic. This approach has been the first solution to achieve real-time rendering of high-resolution, photorealistic novel views for complex dynamic scenes. To facilitate real-world applications, we derive several compact variants that effectively reduce the memory footprint to address its storage bottleneck. Extensive experiments validate the superiority of 4DGS in terms of visual quality and efficiency across a range of dynamic scene-related tasks (e.g., novel view synthesis, 4D generation, scene understanding) and scenarios (e.g., single object, indoor scenes, driving environments, synthetic and real data).
keywords:
Dynamic scenes, Novel view synthesis, Gaussian splatting, Neural video synthesis
1 Introduction
Modeling dynamic scenes and rendering novel views is crucial in computer vision and graphics, which has received increasing attention from both industry and academia because of its potential value in a wide range of AR/VR applications. Whilst recent breakthroughs (e.g., NeRF [1]) have achieved photorealistic rendering for static scenes [2, 3], handling dynamic scenes remains a significant challenge. The temporal dynamics of scenes introduce additional complexity. The central challenge lies in preserving intrinsic correlations and sharing relevant information across different time steps to create a compact representation, while minimizing interference between unrelated spacetime locations.
@@ 1 Introduction
The second group posits that scene dynamics underlie the motion and deformation of a canonical 3D model [9, 10, 11, 12]. These approaches explicitly model scene motion, potentially improving the utilization of spatial and temporal correlations. However, global tracking of object motion involves inherent ambiguities and risks singularities when dealing with complex motions in real-world scenes, such as sudden appearances and disappearances, which limits their capability and scalability in unconstrained environments.
@@
To address these fundamental limitations, in this work we reformulate the modeling of dynamic scenes as a generic, spatio-temporal 4D volume learning problem. This novel perspective not only introduces an explicit representation, but also makes minimal assumptions about how motion is composed, paving the way for a versatile dynamic scene learning framework. Concretely, we propose to use a native 4D primitive 4D Gaussian that models the geometry and motion of a time-varying 3D scene at the same time. The adopted 4D Gaussian parameterized by anisotropic ellipses that can rotate arbitrarily in space and time fits naturally within the intrinsic spatio-temporal manifold. Furthermore, we introduce Spherindrical Harmonics, a generalized form of Spherical Harmonics, to represent the time evolution of appearance in dynamic scenes, along with a dedicated splatting-based differentiable rendering pipeline. Our 4DGS model is the first to support end-to-end training and real-time rendering of high-resolution, photorealistic novel views of complex dynamic scenes.
@@
We make the following contributions: (i) We propose to use a set of anisotropic 4D Gaussians equipped with 4D Spherindrical Harmonics to represent the dynamic 3D scene, which is the first native 4D primitive. (ii) We cast the novel view synthesis of the dynamic scene as to optimize the proposed 4D Gaussian primitives for direct fitting the underlying spatio-temporal volume, and introduce the necessary differentiable rasterization pipeline. (iii) Extensive experiments validate the superiority of 4DGS over previous alternatives in terms of visual quality and efficiency across a variety of dynamic scene-based tasks (e.g., novel view synthesis, content generation, scene understanding) and scenarios (e.g., single object, indoor and outdoor scenes, synthetic and real data).
@@
A preliminary version of this work has been presented [13]. Moving beyond our initial focus on novel view synthesis, here we significantly extend and explore the proposed 4DGS from a versatile dynamic scene representation perspective and manage to address the storage bottleneck. In comparison, we summarize several important improvements: (i) Introducing more compact variants: There is a trade-off between compactness and versatility governed by how much assumption of the scene’s dynamic hold by the representation. We prioritize the latter while leaving room to inject motion priors via regularization. As a result, the size of vanilla 4DGS may increase rapidly with the inherent large 4D volume. However, this is not a fundamental limitation. We introduce several model compression strategies across multiple dimensions, including parameter size, the number of Gaussian points, that substantially alleviate storage demands and reduce memory footprint. (ii) Application extension: We first extend our model to handle large-scale dynamic urban scenes, which present additional challenges such as sparse viewpoints and unbounded environments. We also enhance the ability to create 4D content from monocular video via a generative approach. Finally, we demonstrate the proposed representation can enhance time-consistent scene understanding, providing object segmentation across both space and time. (iii) Extensive evaluation: We conduct comprehensive evaluations of model compression and the newly added applications on more benchmarks.
2 Related work
2.1 Novel view synthesis
@@
Novel view synthesis for static scenes The field of novel view synthesis has received widespread attention initiated by NeRF [1]. Querying a MLP for hundreds of points along each ray, NeRF is significantly limited in training and rendering. Subsequent works attempted to improve the speed [14, 15, 16, 17, 18, 19] and enhance rendering quality [20, 3, 2, 21, 22]. Later, 3D Gaussian Splatting (3DGS) [23] achieves real-time rendering at high resolution thanks to its fast differentiable rasterization pipeline while retaining the merits of volumetric representations. Subsequent works [24, 25] have improved rendering [26, 27, 28, 29, 30] and geometric accuracy [31, 32, 33]. Here, we explore the potential of Gaussian primitives in representing dynamic scenes.
@@
NeRF-based dynamic view synthesis Synthesizing novel views of a dynamic scene in a desired time is a more challenging task. The intricacy lies in capturing the intrinsic correlation across different timesteps. Inspired by NeRF, one research line attempts to learn a 6D plenoptic function represented by a well-tailored structure [4, 6, 7, 5, 8]. However, these methods struggle with the coupling between parameters. An alternative approach explicitly models continuous motion or deformation [9, 10, 11, 12]. Among them, point-based approaches have consistently been deemed promising.
@@
Compression of 3DGS Despite impressive novel view synthesis performance, 3DGS imposes substantial storage requirements due to the associated attributes of a large number of Gaussians. To reduce the memory footprint, [34] uses the hash grid to encode the color and quantifies other parameters through residual vector quantization (RVQ) [35], while utilizing a learned volume mask to eliminate Gaussians with small size. [36] introduces a scale- and resolution-aware redundancy score as a criterion for removing redundant primitives. [37] compresses parameters of part of Gaussians in a sensitivity-aware manner. [38] stores the related Gaussians in the same anchor and dynamically spawns neural Gaussians during rendering. [39] further organizes sparse anchor points using a structured hash grid. We systematically explore these compression approaches with our dynamic scene representations.
@@
Dynamic 3D Gaussians There have been recent efforts [12, 40, 41, 42, 43, 44, 45, 46, 47] in extending 3DGS for dynamic scenes. A straight approach is to incrementally reconstruct the dynamic scene via frame-by-frame training [12]. [40, 41, 42] model the geometry and dynamics of scenes by joint optimizing Gaussians in canonical space and a deformation field. [43] encourages locality and rigidity between points by factorizing the motion of the scene into a few neural trajectories. These representations hold the topological invariance and low-frequency motion prior, thus well-suited for reconstructing dynamic scenes from monocular videos. However, they assume that dynamic scenes are induced by a fixed set of 3D Gaussians and that all scene elements remain visible. Circumventing the need to maintain ambiguous global tracking, our approach can more flexibly deal with complex real-world scenes.
2.2 Applications with 3D/4D scene representations
@@
Driving scene synthesis Efficient simulation of realistic sensor data in diverse driving scenes is crucial for scaling autonomous driving systems. Early works [48, 49] employ traditional graphics techniques but suffer from domain gap of synthesized data and laborious manual modeling. [50] creates realistic novel driving videos by inserting 3D assets according to the depth map.  [51, 52, 53, 54, 55] utilize neural implicit fields to reconstruct the scene from captured driving data, enabling novel view synthesis. Recent works [56, 57, 58, 59] have applied 3DGS to driving scenes, significantly improving fidelity while achieving real-time simulation. Other notable efforts focus on reconstructing scenes without relying on 2D masks or 3D bounding boxes of dynamic objects. Among them, SUDS [60] and EmerNeRF [61] decompose dynamic and static regions by optimizing optical flow, while $S^{3}$Gaussian [62] and PVG [63] achieve this through self-supervision.
@@
3D/4D content generation Generative capabilities of 3D/4D representations have been explored. In 3D generation, DreamFields [64] and DreamFusion [65] utilize NeRF as a base 3D representation. Magic3D [66] incorporates a textured mesh based on the deformable tetrahedral grid in the refinement stage. Also, 3DGS has been applied to 3D generation [67, 68, 69, 70]. For 4D generation, 4D radiance fields [7, 71, 6] have been adopted [72, 71, 73, 74, 75], followed by Gaussian-based representations [40, 41, 13, 45] explored in [76, 77, 78, 79, 80, 81, 82]. We show the advantages of 4DGS in terms of both generation quality and speed in video-to-4D generation task.
@@
3D/4D scene segmentation Following the success of Segment Anything Model (SAM) [83] in 2D image segmentation, recent works [84, 85, 86, 87, 88, 89, 90, 91] extend NeRF and 3DGS for scene understanding and segmentation. SAGA [85] applies multi-granularity segmentation to 3DGS using a scale gate mechanism. SA4D [84] conducts 4D segmentation based on deformed 3DGS with single-granularity. We validate the general efficacy of 4DGS for 4D scene understanding.
3 Method
@@
We propose a generic scene representation, 4D Gaussian splatting (4DGS), for modeling dynamic scenes, as shown in Figure 2. Next, we will delineate each component of 4DGS and corresponding optimization process. In Section 3.1, we review 3D Gaussian splatting [23]. In Section 3.2, we detail how to represent dynamic scenes and synthesize novel views with our native 4D primitives, i.e., 4D Gaussian. Finally, the optimization framework will be introduced in Section 3.3.
3.1 Preliminary: 3D Gaussian splatting
@@
Representation of 3D Gaussians In 3DGS, a scene is represented as a cloud of 3D Gaussians. Each Gaussian has a theoretically infinite scope and its influence on a given spatial position $x\in\mathbb{R}^{3}$ defined by an unnormalized Gaussian function:
@@
Differentiable rasterization via Gaussian splatting In rendering, given a pixel $(u,v)$ in view $\mathcal{I}$ with extrinsic matrix $E$ and intrinsic matrix $K$, its color $\mathcal{I}(u,v)$ can be computed by blending visible 3D Gaussians that have been sorted according to their depth, as described below:
@@
3.2 4D Gaussian for dynamic scenes
Reformulation To enable Gaussian splatting for modeling dynamic scenes, reformulation is necessary. In dynamic scenes, a pixel under view $\mathcal{I}$ can no longer be indexed solely by a pair of spatial coordinates $(u,v)$ in the image plane; But an additional timestamp $t$ comes into play and intervenes. Formally this is formulated by extending equation (2) as:
$\mathcal{I}(u,v,t)=$ $\sum^{N}_{i=1}(p_{i}(u,v,t)\alpha_{i}c_{i}(d)$ (5)
$\prod^{i-1}_{j=1}(1-p_{j}(u,v,t)\alpha_{j})).$
@@
$\mathcal{I}(u,v,t)=$ $\sum^{N}_{i=1}(p_{i}(t)p_{i}(u,v|t)\alpha_{i}c_{i}(d)$ (6)
$\prod^{i-1}_{j=1}(1-p_{j}(t)p_{j}(u,v|t)\alpha_{j})).$
@@
Subsequently, it comes to the question of how to represent a 4D Gaussian. An intuitive solution is that we adopt a distinct perspective for space and time. Consider that $(x,y,z)$ and $t$ are independent of each other, i.e., $p_{i}(x,y,z|t)=p_{i}(x,y,z)$, then equation (6) can be implemented by adding an extra 1D Gaussian $p_{i}(t)$ into the original 3D Gaussian. This design can be viewed as imbuing a 3D Gaussian with temporal extension, or weighting down its opacity when the rendering timestep is away from the expectation of $p_{i}(t)$. However, our experiment shows that while this approach can achieve a reasonable fitting of 4D manifold, it is difficult to capture the underlying motion of the scene (see Section 8.1.3 for details).
Representation of 4D Gaussian To address the mentioned challenge, we suggest treating the time and space dimensions equally by formulating a coherent integrated 4D Gaussian model. Similar to [23], we parameterize its covariance matrix $\Sigma$ as the configuration of a 4D ellipsoid for easing model optimization:
@@
Proof of unnormalized Gaussian property The previous formulation treats the unnormalized Gaussian defined in equation (1) as a specialized probability distribution and assumes it can be factorized into the product of the unnormalized Gaussian functions of its conditional and marginal distributions. In this paragraph, we provide a concise proof for this property.
@@
4D spherindrical harmonics The view-dependent color $c_{i}(d)$ in equation (6) is represented by a series of SH coefficients in the original 3DGS. To more faithfully model the dynamic scenes of the real world, we need to enable appearance variation with both varying viewpoints and evolving time.
@@
where $Y^{m}_{l}$ is the 3D spherical harmonics. The index $l\geq 0$ denotes its degree, and $m$ is the order satisfying $l\leq m\leq l$. The index $n$ is the order of the Fourier series. The 4D spherindrical harmonics form an orthonormal basis in the spherindrical coordinate system.
3.3 Training